OpenStudy (anonymous):

what is the period of the function:

OpenStudy (anonymous):

\[f(\theta)=-2\cos 3\theta \]

hartnn (hartnn):

the standard form is : \(f(\theta)=A\cos (\omega t+\phi)=A\cos(2\pi t/T+\phi)\) where T is the time period. here, we get , \(\omega=2\pi/T = 3\) can you find t from here ?

OpenStudy (anonymous):

you lost me

hartnn (hartnn):

i just compared cos \(3\theta\) with \(\cos \omega t\) to get w=3 and time period T can be calculated from w= 2pi/T so, \(2\pi /T=3\) can't you find T from here ?

OpenStudy (anonymous):

i dont think ive learned that yet

hartnn (hartnn):

now you have , its really easy to get T from that equation....try once.

OpenStudy (anonymous):

the only thing is i think my teacher wanted us to review idk if she'd be able to give me credit for that

OpenStudy (anonymous):

omg i figured it out thank you though its all out of the unit circle!!

OpenStudy (anonymous):

its the formula 2pi/b

OpenStudy (anonymous):

... you used t and i got confused thank you so so much!

hartnn (hartnn):

sorry for confusing you, i learned it that way, and yes, its 2pi/b. welcome :)

OpenStudy (anonymous):

lol guess thats what one letter can do . but thank you so much any way